# GBR: Generative Bundle Refinement for High-fidelity Gaussian Splatting and Meshing

### v1

@@ I Introduction
However, existing Gaussian splatting methods still require dense-view inputs (approximately 100 views) to achieve a good performance. When working with sparse-view inputs ($<10$ views), challenges such as background collapse and the presence of excessive floaters frequently arise, leading to inaccuracies in geometry reconstruction and degraded quality in novel view synthesis (NVS). For instance, the original Gaussian splatting pipeline[7] utilizes structure-from-motion (SfM) methods like COLMAP [8] to get the initial view poses and point cloud. Unfortunately, COLMAP struggles in sparse-view inputs, as it fails to generate a sufficiently dense and complete point cloud for the initialization of Gaussian primitives, thereby limiting the geometric accuracy and mesh fidelity. Recent methods [9, 10] also explored COLMAP-free initialization methods, but the dense-view inputs are still necessary.
@@ II Related Work
Building on the success of NVS in 3D representation, methods like NeRF and 3DGS have also been adapted for surface reconstruction tasks. NeRF-based approaches for surface modeling have incorporated occupancy fields [23] and signed distance fields [24, 25] to capture surface geometry. Although effective, the reliance of NeRF on MLP layers can be a bottleneck for both inference speed and flexibility in representation. To mitigate these limitations, recent approaches are moving away from MLP-heavy architectures, instead using discrete structures like points [26, 27] and voxels [28, 29, 30] to represent scene information more efficiently.
SuGaR [31] introduces a technique for extracting meshes from 3DGS by employing regularization terms to better align Gaussians with scene surfaces. This method samples 3D points from the Gaussian density field and applies Poisson surface reconstruction for meshing. While this improves geometric fidelity, the irregular shapes of 3D Gaussians can complicate smooth surface modeling, and the unordered structure of Gaussians makes the image reconstruction loss vulnerable to overfitting, potentially leading to incomplete geometry and surface misalignment issues.
To enhance multi-view consistency, methods such as 2DGS [32] compress the 3D volume into planar Gaussian disks, while GOF [33] uses a Gaussian opacity field to extract geometry through level sets directly. PGSR [34] transforms Gaussian shapes into planar forms, optimizing them for realistic surface representation and simplifying parameter calculations, such as normals and distance measures. Nonetheless, achieving consistent geometry across multiple views remains a challenge in 3DGS approaches.
@@ II-C Unconstrained 3D Representations
Several methods introduce regularization to support sparse-view scenarios. Techniques like RegNeRF [35], DGRNeRF [36], GeoRGS [37], and SparseNeRF [38] use depth priors to improve geometric accuracy, while DietNeRF [39] and ReconFusion [40] incorporate guidance from models such as CLIP [41] and diffusion networks [42]. Approaches like PixelNeRF [43] and FreeNeRF [44] utilize pre-training for few-shot NVS, while SparseGS [13], FSGS [14] and DNGaussian [45] employs monocular depth estimators for scenarios with limited views. However, these methods typically assume known camera poses derived from dense image captures.
To bypass the requirement for pre-determined camera poses, recent pose-free methods jointly optimize camera parameters with the 3D model using uncalibrated images. Strategies like NeRFmm [46] and BARF [47] adopt a coarse-to-fine approach for pose encoding. Innovations like Nope-NeRF [10], LuNeRF [48], and CF-3DGS [9] use depth priors to reduce the number of views required. However, they often assume densely captured video sequences, which can lead to optimization times similar to traditional SfM workflows.
Methods like InstantSplat [11] and LM-Gaussian [49] leverage DUSt3R [15] as an initializer, enabling estimation of camera parameters and a relatively dense point cloud. However, the noise inherent in DUSt3R’s initialization significantly affects geometric precision, allowing only visually appealing novel view synthesis rather than accurate geometry reconstruction.
@@ IV-A2 Bundle adjustment refinement with dual filtering
Finally, given the 2D matching points $\mathcal{M}$, intrinsic camera matrix $\mathcal{K}$, extrinsic camera matrix $\mathcal{T}$, and the initial point cloud ${X}$, the bundle adjustment is formulate as:
@@ IV-B1 Diffusion-based iterative depth refinement
  • Step 1: Initial depth map projection: We project the 3D point cloud into 2D depth maps. Limited by the processing resolution of the DUSt3R network, the initial depth map $D^{0}$ usually lacks fine details.
  • Step 2: Latent encoding: The depth map and the corresponding RGB image are input to separate latent encoders, producing latent representations $z^{(D)}$ (depth) and $z^{(I)}$ (RGB).
  • Step 3: Diffusion-based refinement: The latent depth representation $z^{(D)}$ undergoes an iterative refinement process guided by a diffusion model, operating entirely in the latent code space. Starting from the initial latent depth representation $z_{0}^{(D)}$, the process consists of two main stages: (1) In the forward process, Gaussian noise is gradually added to $z_{0}^{(D)}$ over $T$ steps to generate a sequence of noisy latent representations $\{z_{1}^{(D)},z_{2}^{(D)},\dots,z_{T}^{(D)}\}$. At each step $t$, this process is given by:
@@
  • Step 4: Latent decoding: The refined depth latent is decoded into a high-resolution depth map, effectively addressing the resolution limitations of the DUSt3R network.
@@ IV-B2 Scale-consistent depth integration
$\epsilon_{i}$ =$=\begin{cases}\epsilon_{\text{edge}},&\text{if }|\nabla D^{0}_{i}|\geq\tau_{e}\\ \epsilon_{\text{smooth}},&\text{if }|\nabla D^{0}_{i}|<\tau_{e}\end{cases},$
@@ IV-C5 Photometric loss
We set $\lambda_{1}=0.005$,$\lambda_{2}=0.005$,$\lambda_{3}=0.1$,and $\lambda_{4}=0.1$. Upon completing the training of the Gaussian primitives, we utilize the TSDF fusion method [53] to merge the depth maps from all views, resulting in the final mesh.
@@ V-A Dataset and sparse view selection
Figure 6: Qualitative results using 6 input views on TNT dataset. The figure shows surface reconstruction quality using normal maps, with sky regions masked out. Our method generates smooth, detail-rich geometry, outperforming others that struggle with noise and fail to capture fine details. Additionally, our approach reconstructs both foreground and background geometry, unlike other methods restricted to foreground objects.
Figure 7: Qualitative NVS results using 6 input views on the MIP dataset. The figure demonstrates the synthesized RGB images and normal maps of different methods. Our approach renders novel views with sharper details and fewer artifacts.
Similar to previous algorithms, we selected the DTU [54], TNT [55], and MIP [18] datasets for testing our method. The DTU and TNT datasets are primarily used to evaluate geometric reconstruction performance, while the MIP dataset is employed to assess the algorithm’s ability to generate novel views and to qualitatively showcase the mesh reconstruction results. For the DTU dataset, we selected 4 views for training and testing. This sparse-view configuration presents a significant challenge. For the larger scene datasets (TNT and MIP), we used 6 views as training data and an additional 6 neighboring views for testing. For the DTU and TNT datasets, we sampled the ground truth 3D geometry and only preserved those points visible from two or more views. For the TNT dataset testing, irrelevant point clouds, such as sky regions and background vegetation, were excluded to improve the quality of the results.
@@ V-B Evaluation criteria
For the novel view synthesis task, we employed three widely used image evaluation metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS). For surface quality evaluation, we assessed performance using the $F_{1}$ score (TNT) and Chamfer Distance (DTU). Additionally, we report the Absolute Trajectory Error (ATE) to evaluate the effectiveness of the camera parameters optimization. The real camera poses provided by the dataset are used as the ground truth.
@@ V-D Baselines
We combine DUSt3R [15] with other Gaussian splatting-based surface reconstruction algorithms as the primary baseline. For the optimization of DUSt3R, we employ the method of InstantSplat [11] to average the intrinsic parameters, which can improve the accuracy of the generated 3D point cloud. Additionally, we also conducted experiments using COLMAP combined with the Gaussian splatting-based surface reconstruction algorithms. However, this approach did not yield satisfactory results on the TNT dataset and is not included in our evaluations. For surface reconstruction algorithms, we select 2DGS [32], GOF [33], and PGSR [34]. We excluded SuGaR [31] from our evaluation due to its relatively high failure rate.
TABLE I: Quantitative comparison of methods on DTU data (Chamfer distance mm $\downarrow$). Red cells indicate the best performance, and light red cells indicate the second-best performance for each metric. ‘-‘ indicates experiment failure.
@@
2DGS+DUSt3R 0.846 4.127 4.197 4.594 5.133 5.457 5.030 1.020 1.453 4.985 4.530 0.727 4.560 5.084 1.670 3.561
PGSR+DUSt3R 0.948 4.740 3.836 4.371 1.623 2.002 4.628 0.641 1.349 4.696 0.669 0.726 4.334 4.379 6.451 3.026
GOF+DUSt3R 1.181 3.466 4.211 4.537 4.720 5.263 4.495 0.863 6.300 6.084 4.373 0.727 3.568 4.604 5.709 4.007
@@
V-E DTU dataset
Table I presents the quantitative results of the geometric evaluations for the DTU dataset. As discussed earlier, network-based geometry estimation methods, such as DUSt3R, suffer from low accuracy. While the conventional COLMAP method can produce highly accurate results, it is less robust for sparse-view inputs. For instance, COLMAP fails to reconstruct the 100 data with only 4-view inputs. While our method combines the advantages of these two methods, significantly outperforming the baseline methods in most cases. Only in a few specific scenarios, such as scenes 55, 93, and 114, where COLMAP can provide accurate extrinsic parameters, the baseline methods show marginal competitiveness with our method.
V-F TNT dataset
Compared to the DTU dataset, the Tank and Temples (TNT) dataset collects outdoor real-world data, which presents more significant challenges. The numerical evaluation results in Table II demonstrate that our method holds a clear advantage across all scenes. Fig. 6 shows the rendered normal maps. Our normal map is more complete and cleaner than the baseline methods, with better object details. For example, in our results, the eyes, nose, and fingers of the sculpture are successfully reconstructed, while the other methods fail to generate reasonable meshes and normal maps.
@@ V-F TNT dataset
TABLE II: Quantitative result of $F_{1}$ $\uparrow$ on Tanks and Temples dataset.
V-G MIP dataset
We use the MIP dataset to test the performance of novel view synthesis. In addition to the baseline methods, we also evaluated two pose-free methods: InstantSplat and COLMAP-free 3D Gaussian splatting (CF-3DGS) [9]. InstantSplat uses DUSt3R for initialization but without extensive optimization, simply setting the focal length to an average value. CF-3DGS processes the input frames in a sequential manner and progressively grows the 3D Gaussians set by taking one input frame at a time.
@@ V-G MIP dataset
TABLE III: Quantitative comparison of NVS task on MIP dataset. PSNR and SSIM are higher the better, while LPIPS is lower the better.
V-H Ablations
Table IV presents the results of our ablation study, which evaluates the contribution of the neural-BA module, depth refinement module, depth loss, geometric consistency loss, and pseudo view loss. The Barn scene from the TNT dataset is used for testing. Each row corresponds to a different configuration of the system, with one key component excluded. Precision, recall, and $F_{1}$ score are used to show the impact of all the components. The results effectively highlight the critical role of all the components in the final performance.
@@ V-H Ablations
TABLE IV: Quantitative result of $F_{1}$ $\uparrow$ on Barn scene of TNT dataset.

### v2

# GBR: Generative Bundle Refinement for High-fidelity Gaussian Splatting with Enhanced Mesh Reconstruction

@@ I Introduction
However, existing Gaussian splatting methods still require dense-view inputs (approximately 100 views) to achieve a good performance. When working with sparse-view inputs (<10<10 views), challenges such as background collapse and the presence of excessive floaters frequently arise, leading to inaccuracies in geometry reconstruction and degraded quality in novel view synthesis (NVS). For instance, the original Gaussian splatting pipeline[7] utilizes structure-from-motion (SfM) methods like COLMAP [8] to get the initial view poses and point cloud. Unfortunately, COLMAP struggles in sparse-view inputs, as it fails to generate a sufficiently dense and complete point cloud for the initialization of Gaussian primitives, thereby limiting the geometric accuracy and mesh fidelity. Recent methods [9, 10] also explored COLMAP-free initialization methods, but the dense-view inputs are still necessary.
@@ II Related Work
Surface reconstruction is a classic problem in 3D vision, aimed at reconstructing high-quality 3D assets from visual data. Methods like MVSNet [23] utilize multi-view stereo to generate depth maps and point clouds, enabling effective surface reconstruction through traditional geometric representations. SurRF [24] introduces a surface radiance field defined on a continuous 2D surface, achieving compact representation while maintaining realistic texture and high geometric resolution. Building on the success of NVS in 3D representation, methods like NeRF and 3DGS have also been adapted for surface reconstruction tasks. NeRF-based approaches for surface modeling have incorporated occupancy fields [25] and signed distance fields [26, 27] to capture surface geometry. Although effective, the reliance of NeRF on MLP layers can be a bottleneck for both inference speed and flexibility in representation. To mitigate these limitations, recent approaches are moving away from MLP-heavy architectures, instead using discrete structures like points [28, 29] and voxels [30, 31, 32] to represent scene information more efficiently.
SuGaR [33] introduces a technique for extracting meshes from 3DGS by employing regularization terms to better align Gaussians with scene surfaces. This method samples 3D points from the Gaussian density field and applies Poisson surface reconstruction for meshing. While this improves geometric fidelity, the irregular shapes of 3D Gaussians can complicate smooth surface modeling, and the unordered structure of Gaussians makes the image reconstruction loss vulnerable to overfitting, potentially leading to incomplete geometry and surface misalignment issues. To enhance multi-view consistency, methods such as 2DGS [34] compress the 3D volume into planar Gaussian disks, while GOF [35] uses a Gaussian opacity field to extract geometry through level sets directly. PGSR [36] transforms Gaussian shapes into planar forms, optimizing them for realistic surface representation and simplifying parameter calculations, such as normals and distance measures. Nonetheless, achieving consistent geometry across multiple views remains a challenge in 3DGS approaches.
@@ II-C Unconstrained 3D Representations
Several methods have been proposed to tackle the challenge of novel view synthesis (NVS) from sparse-input views. NeRF-based approaches often introduce regularization to compensate for missing visual information and improve reconstruction quality. RegNeRF [37] enforces geometric consistency through ray sampling annealing and normalizing flow, while DGRNeRF [38] integrates point cloud priors to refine depth accuracy, making it more robust to sparse inputs. SparseNeRF [39] further enhances depth estimation by incorporating local depth ranking and spatial continuity constraints, reducing reconstruction artifacts in under-constrained regions. FreeNeRF [40] enhances sparse-view synthesis by applying frequency regularization, which suppresses overfitting to incomplete observations and promotes more stable reconstructions.
Beyond regularization, some methods incorporate external guidance to infer missing details in sparse-view scenarios. DietNeRF [41] utilizes CLIP-based semantic consistency to enforce high-level feature alignment across different viewpoints, thereby improving scene coherence. ReconFusion [42] employs a diffusion prior to refine both geometric and textural fidelity, helping NeRF reconstruct plausible details even from a limited number of input images. Other approaches leverage feature information to enhance NeRF’s adaptability to sparse settings. PixelNeRF [43] uses convolutional feature encodings to learn scene priors, enabling generalization across multiple scenes with minimal input views. MVSNeRF [44] employs plane-swept cost volumes to provide geometry-aware feature information, allowing for efficient radiance field reconstruction from just three views while ensuring strong scene generalization ability.
Beyond NeRF, 3D Gaussian Splatting-based methods have recently received attention due to their efficiency and rendering quality in sparse-view settings. GeoRGS [45] regularizes the growth of 3D Gaussian points to mitigate early-stage errors, which is crucial in sparse-input settings where incorrect initial geometry can lead to severe overfitting. By guiding Gaussian expansion with seed patches and enforcing depth similarity, GeoRGS improves scene consistency and reduces artifacts. SparseGS [13] incorporates depth priors and generative constraints to prevent common artifacts caused by sparse-input views such as background collapse. FSGS [14] enables real-time, photo-realistic rendering by utilizing Gaussian unpooling and monocular depth estimation, requiring as few as three input views. DNGaussian [46] further enhances efficiency by introducing an adaptive depth regularization strategy that targetedly constrains Gaussian parameters rather than applying global depth supervision. This approach improves geometry learning while reducing artifacts, achieving a 25× faster training speed and over 3000× faster rendering compared to previous methods.
Despite these advancements, most methods still assume known camera poses, often derived from dense captures or precomputed estimates. In unconstrained settings, where pose uncertainty is prevalent, further research is needed to enhance robustness against incomplete or ambiguous camera parameters. Recently, pose-free methods are proposed to jointly optimize camera parameters and the 3D model. Strategies like NeRFmm [47] and BARF [48] adopt different approaches for pose encoding, enabling robust pose estimation and novel view synthesis. NeRFmm treats camera parameters as learnable variables optimized through photometric reconstruction, while BARF formulates the problem as joint scene representation learning and camera registration, leveraging coarse-to-fine positional encoding to mitigate misalignment during training. Nope-NeRF [10], LuNeRF [49], and CF-3DGS [9] utilize depth priors from monocular depth estimation to iteratively refine camera poses. By optimizing poses frame by frame with depth constraints, these methods effectively refine camera poses without relying on pre-computed initialization, making pose-free training feasible. However, these methods typically rely on densely captured video sequences as input, making them less effective in handling sparse-view scenarios. Methods like InstantSplat [11] and LM-Gaussian [12] utilize DUSt3R [15] as an initializer to enable the the estimation of camera parameters and a relatively dense point cloud for training 3D Gaussian Splatting (3DGS). Similarly, ViewCrafter [50] and ReconX [51] leverage the dense initialization generated by DUSt3R and employ video diffusion techniques to synthesize high-quality multi-view images. However, the noise inherent in DUSt3R’s initialization significantly affects geometric precision, allowing only visually appealing novel view synthesis rather than accurate geometry reconstruction.
In summary, achieving high-fidelity novel view synthesis and accurate 3D reconstruction in unconstrained scenarios, where camera poses are unknown and viewpoints are sparse, remains a significant challenge. The limited observations lead to ambiguities in both geometric estimation and 3D representation optimization, requiring effective solutions to improve reconstruction quality.
@@ IV-A2 Bundle adjustment refinement with dual filtering
Finally, given the 2D matching points $\mathcal{M}$, intrinsic camera matrix $\mathcal{K}$, extrinsic camera matrix $\mathcal{T}$, and the initial point cloud X{X}, the bundle adjustment is formulate as:
@@ IV-B1 Diffusion-based iterative depth refinement
  • Step 1: Initial depth map projection. We project the 3D point cloud into 2D depth maps. Limited by the processing resolution of the DUSt3R network, the initial depth map $D^{0}$ usually lacks fine details.
  • Step 2: Latent encoding. The depth map and the corresponding RGB image are input to separate latent encoders, producing latent representations $z^{(D)}$ (depth) and $z^{(I)}$ (RGB).
  • Step 3: Diffusion-based refinement. The latent depth representation $z^{(D)}$ undergoes an iterative refinement process guided by a diffusion model, operating entirely in the latent code space. Starting from the initial latent depth representation $z_{0}^{(D)}$, the process consists of two main stages: (1) In the forward process, Gaussian noise is gradually added to $z_{0}^{(D)}$ over $T$ steps to generate a sequence of noisy latent representations $\{z_{1}^{(D)},z_{2}^{(D)},\dots,z_{T}^{(D)}\}$. At each step $t$, this process is given by:
@@
  • Step 4: Latent decoding. The refined depth latent is decoded into a high-resolution depth map, effectively addressing the resolution limitations of the DUSt3R network.
@@ IV-B2 Scale-consistent depth integration
[…]
@@ IV-C5 Photometric loss
We set $\lambda_{1}=0.005$$\lambda_{2}=0.005$$\lambda_{3}=0.1$,and $\lambda_{4}=0.1$. Upon completing the training of the Gaussian primitives, we utilize the TSDF fusion method [55] to merge the depth maps from all views, resulting in the final mesh.
@@ V-A Dataset and sparse view selection
Figure 6: Qualitative results using 6 input views on Tanks & Temples dataset. The figure shows surface reconstruction quality using normal maps, with sky regions masked out. Our method generates smooth, detail-rich geometry, outperforming others that struggle with noise and fail to capture fine details. Additionally, our approach reconstructs both foreground and background geometry, unlike other methods restricted to foreground objects.
Figure 7: Qualitative NVS results using 6 input views on the Mip-NeRF 360 dataset. The figure demonstrates the synthesized RGB images and normal maps of different methods. Our approach renders novel views with sharper details and fewer artifacts.
Similar to previous algorithms, we selected the DTU [56], Tanks & Temples [57], and Mip-NeRF 360 [18] datasets for testing our method. The DTU and TNT datasets are primarily used to evaluate geometric reconstruction performance, while the Mip-NeRF 360 dataset is employed to assess the algorithm’s ability to generate novel views and to qualitatively showcase the mesh reconstruction results. For the DTU dataset, we selected 4 views for training and testing. This sparse-view configuration presents a significant challenge. For the larger scene datasets (Tanks & Temples and Mip-NeRF 360), we used 6 views as training data and an additional 6 neighboring views for testing. For the DTU and Tanks & Temples datasets, we sampled the ground truth 3D geometry and only preserved those points visible from two or more views. For the Tanks & Temples dataset testing, irrelevant point clouds, such as sky regions and background vegetation, were excluded to improve the quality of the results.
@@ V-B Evaluation criteria
For the novel view synthesis task, we employed three widely used image evaluation metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS). For surface quality evaluation, we assessed performance using the $F_{1}$ score (Tanks & Temples) and Chamfer Distance (DTU). Additionally, we report the Absolute Trajectory Error (ATE) to evaluate the effectiveness of the camera parameters optimization. The real camera poses provided by the dataset are used as the ground truth.
@@ V-D Baselines
We combine DUSt3R [15] with other Gaussian splatting-based surface reconstruction algorithms as the primary baseline. For the optimization of DUSt3R, we employ the method of InstantSplat [11] to average the intrinsic parameters, which can improve the accuracy of the generated 3D point cloud. Additionally, we also conducted experiments using COLMAP combined with the Gaussian splatting-based surface reconstruction algorithms. However, this approach did not yield satisfactory results on the Tanks & Temples dataset and is not included in our evaluations. For surface reconstruction algorithms, we select 2DGS [34], GOF [35], and PGSR [36]. We excluded SuGaR [33] from our evaluation due to its relatively high failure rate.
@@
2DGS+DUSt3R 0.833 1.475 1.759 0.995 1.779 1.598 1.460 1.649 1.805 1.986 1.546 1.943 1.296 1.419 1.289 1.522
PGSR+DUSt3R 0.950 1.685 1.322 1.028 1.545 1.823 1.366 1.921 1.793 1.386 1.379 2.387 1.154 1.186 1.140 1.471
GOF+DUSt3R 1.098 1.825 2.130 1.196 1.766 2.077 1.608 1.880 1.997 1.935 1.862 1.892 1.347 2.072 2.721 1.827
@@
TABLE I: Quantitative comparison of methods on DTU data (Chamfer distance $m$ $\downarrow$). Red cells indicate the best performance, and light red cells indicate the second-best performance for each metric. ‘-‘ indicates experiment failure.
@@
[…]
@@
TABLE III: Quantitative comparison of NVS task on Mip-NeRF 360 dataset. PSNR and SSIM are higher the better, while LPIPS is lower the better.
Method 3 views 4 views 5 views 6 views
2DGS+DUSt3R 1.589 1.522 1.454 1.438
PGSR+DUSt3R 1.478 1.471 1.464 1.358
Ours 0.543 0.526 0.520 0.521
TABLE IV: Quantitative comparison of different methods on the DTU dataset (Chamfer distance $m$ $\downarrow$) under varying numbers of input views.
@@
[…]
V-E Geometric Evaluation on the Synthetic Dataset
Table I presents the quantitative results of the geometric evaluations for the synthetic DTU dataset. As discussed earlier, network-based geometry estimation methods, such as DUSt3R, suffer from low accuracy. While the conventional COLMAP method can produce highly accurate results, it is less robust for sparse-view inputs. For instance, COLMAP fails to reconstruct the 100 data with only 4-view inputs. While our method combines the advantages of these two methods, significantly outperforming the baseline methods in most cases. Only in a few specific scenarios, such as scenes 55, 93, and 114, where COLMAP can provide accurate extrinsic parameters, the baseline methods show marginal competitiveness with our method.
V-F Geometric Evaluation on Real-world Large-scale Scenes
Compared to the DTU dataset, the Tanks & Temples dataset collects outdoor real-world data, which presents more significant challenges. The numerical evaluation results in Table II demonstrate that our method holds a clear advantage across all scenes. Fig. 6 shows the rendered normal maps. Our normal map is more complete and cleaner than the baseline methods, with better object details. For example, in our results, the eyes, nose, and fingers of the sculpture are successfully reconstructed, while the other methods fail to generate reasonable meshes and normal maps.
V-G Evaluation of Novel View Synthesis and Camera Pose Estimation
We use the Mip-NeRF 360 dataset to test the performance of novel view synthesis and pose estimation. In addition to the baseline methods, we also evaluated two pose-free methods: InstantSplat and COLMAP-free 3D Gaussian splatting (CF-3DGS) [9]. InstantSplat uses DUSt3R for initialization but without extensive optimization, simply setting the focal length to an average value. CF-3DGS processes the input frames in a sequential manner and progressively grows the 3D Gaussians set by taking one input frame at a time.
Figure 8: Qualitative comparison of feature extraction and matching results. Compared to traditional feature matching methods, our approach effectively handles texture-less regions, achieving a more uniform distribution of keypoints across the scene and improving correspondence reliability.
V-H Robustness Testing
V-H1 Robustness of Feature Point Matching
Unlike traditional methods that rely on local image content features, which often struggle in texture-less regions, our algorithm utilizes the powerful feature extraction capabilities of transformers. As shown in Fig. 8, by emphasizing global image-level representations over localized details, our method effectively handles texture-less regions, such as floors, leaves, fruit skins, as well as other large and uniform surfaces. It ensures a more uniform distribution of keypoint correspondences across the scene, enabling more reliable and accurate matching in these challenging areas.
V-H2 Robustness Across Varying Numbers of Views
To further demonstrate the impact of view count on algorithm performance, we supplemented our analysis on the DTU dataset with an additional set of experiments. Specifically, we evaluated the performance under three-view settings for increased difficulty, the existing four-view setting, five-view setting and a six-view setting comparable to the configuration used for Tanks & Temples. The experimental results in Table IV show that our algorithm consistently outperforms others across varying numbers of views, demonstrating its robustness to sparse-view conditions.
V-I Ablations
Table V presents the results of our ablation study, which evaluates the contribution of the neural-BA module, depth refinement module, depth loss, geometric consistency loss, and pseudo view loss. The Barn scene from the Tanks & Temples dataset is used for testing. Each row corresponds to a different configuration of the system, with one key component excluded. Precision, recall, and $F_{1}$ score are used to show the impact of all the components. The results effectively highlight the critical role of all the components in the final performance.